OpenStudy (anonymous):

In ABC, a=4k cm, b=5k cm and c=6k cm. If the area of ABC is 15√7 cm2, find the value of k.

OpenStudy (anonymous):

15sqrt7 = 39.68

OpenStudy (anonymous):

is a, b and c the sides of triangle?

OpenStudy (anonymous):

k value is in cm also, to get cm^2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do we multiply ak * bk * ck to equal 15sqrt7?

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

na just trying to see if u have any further info? i am trying to figure out what to do, to arrive at answer, obviously there is only one value for k

OpenStudy (anonymous):

a,b,c are the sides. nothing else

OpenStudy (anonymous):

ok

OpenStudy (ajprincess):

do u have a diagram? @kryton1212

OpenStudy (anonymous):

no diagram

OpenStudy (ajprincess):

oh k

OpenStudy (ajprincess):

 Area of the triangle=(1/2)*a*h=(1/2)*a*bsinC